# Know When to Abstain: Optimal Selective Classification with Likelihood Ratios

### v1

@@ Appendix A Proofs
is strictly monotone increasing. This can be seen by looking at the first derivative
@@
Since it is positive everywhere in the domain, $h$ is strictly increasing on $[0,\infty)$. Thus we have shown that MSP is a monotone transform of the likelihood ratio $p_{c}/p_{w}$, and Corollary 1 tells us that it is Neyman–Pearson optimal.
@@
where $L=1-d^{(1)}-d^{(2)}\geq 0$. In the binary classification case $L=0$, thus $s_{\text{RLog}}({\bm{x}})=\log\frac{p_{c}({\bm{x}})}{p_{w}({\bm{x}})}$ exactly. In the multiclass case, the $1+L/d^{(2)}$ factor varies across samples and may alter the ordering of scores. If we assume that most of the probability mass is in the top logits as stated in the theorem, which is empirically supported by high top-5 classification accuracies in prior works [45, 46], then $L\approx 0$ and $s_{\text{RLog}}({\bm{x}})\approx\log\frac{p_{c}({\bm{x}})}{p_{w}({\bm{x}})}$. Since logarithm is strictly monotonically increasing, RLog is Neyman–Pearson optimal by Corollary 1. ∎
@@
The empirical likelihood of the KNN density estimator [47, 48] is given by
@@
where $k\geq 2$, $u_{k}({\bm{z}})$ and $v_{k}({\bm{z}})$ are the Euclidean distances from ${\bm{z}}$ to its $k$-th nearest neighbor from $A_{c}$ and $A_{w}$ and $V_{d}$ is the unit-ball volume in $\mathbb{R}^{d}$. A classic result of non-parametric nearest neighbor density estimation [49] states that as $k\rightarrow\infty$ but $k/N_{c}\rightarrow 0$, $k/N_{w}\rightarrow 0$, then $\hat{p}_{c}({\bm{z}})\rightarrow p_{c}({\bm{z}})$ and $\hat{p}_{w}({\bm{z}})\rightarrow p_{w}({\bm{z}})$ for every ${\bm{z}}$. In other words, the empirical KNN density estimator converges to the true density under the stated asymptotic conditions.
@@
$s_{\Delta\text{-KNN}}({\bm{z}})\triangleq-\log u_{k}({\bm{z}})+\log v_{k}({\bm{z}})=\frac{1}{d}\log\frac{\hat{p}_{c}({\bm{z}})}{\hat{p}_{w}({\bm{z}})}-\frac{1}{d}\log\frac{N_{w}}{N_{c}}.$ (12)
@@
Since each score is already a strictly monotone transform of $L({\bm{x}})$, we are free to re-express the scores in any other convenient monotone scale without affecting relative ordering and thus Neyman–Pearson optimality. Therefore, without loss of generality, we will let $s_{i}({\bm{x}})=\log\frac{p_{c}^{(i)}({\bm{x}})}{p_{w}^{(i)}({\bm{x}})},i=1,2$, which are identical to the original scores in terms of sample acceptance and rejection patterns.
Given $\lambda\in\mathbb{R}$,
@@
When is $t({\bm{x}})$ Neyman–Pearson optimal? Recall $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ represent the hypotheses that the classifier will make a correct and wrong prediction respectively. If we assume that the density of $\mathcal{H}_{0}$ takes the form of a tilted likelihood $p_{c}^{(1)}(p_{c}^{(2)})^{\lambda}/Z_{c}$, where $Z_{c}$ is a normalizing constant, and vice-versa for $\mathcal{H}_{1}$, then $t({\bm{x}})$ is Neyman–Pearson optimal by Corollary 1.
Appendix B Discussion on Average Top $k$ $\Delta$-KNN Modification
@@
Note that $U_{i}\in[0,1]$ for all $i$. Since the joint distribution of $U_{i}$ depends only on $k$, we know from the $i$-th order statistics of $k$ i.i.d. $\mathrm{Uniform}(0,1)$ variables that each $U_{i}$ is Beta-distributed:
[…]
With some algebra, $u_{i}({\bm{z}})=\frac{1}{d}\log U_{i}+\log u_{k}({\bm{z}})$. Then we can show that
@@
as $k\to\infty$ as it is a sum of $k$ i.i.d. $\mathrm{Uniform}(0,1)$ random variables.
In other words, in the asymptotic limit the average log-distance is a monotone transform of the log-distance itself. By substituting Eq. 20 back into Eq. 11 for distances to both correct and wrong sets, we see that the modified $\Delta$-KNN formulation is still a monotone transform of the likelihood ratio $p_{c}/p_{w}$, thus suggesting Neyman-Pearson optimality under Corollary 1.
Appendix C Algorithm Pseudocode for Proposed Scores
@@
3:Initialize $\mathcal{A}_{c}^{(i)}\leftarrow\emptyset$ and $\mathcal{A}_{w}^{(i)}\leftarrow\emptyset$ for $i=1$ to $K$triangleright Class-wise correct and incorrect feature sets
@@
3:Initialize $\mathcal{A}_{c}\leftarrow\emptyset$, $\mathcal{A}_{w}\leftarrow\emptyset$triangleright Global sets of correct and incorrect features
@@
Appendix D Experimental Details
@@
Image Experiments.
This section outlines the models and datasets used in the vision experiments in Section 5.1. For classifiers, we use the ViT-H/14 variant for DFN CLIP and the “Giant” variant of EVA, both with patch size 14. Pretrained weights are obtained from the OpenCLIP [^1] https://github.com/mlfoundations/open_clip [50] and timm [^2] https://github.com/huggingface/pytorch-image-models libraries, respectively.
For logit-based scores with CLIP, we are required to construct logits over class concepts by taking the dot product between the text embedding $T_{\theta}({\bm{c}})$ of class concepts ${\bm{c}}$ and the image embedding, $\phi({\bm{x}})$, where $T_{\theta}$ is the text encoder of CLIP. Given a class label $y$, we construct the class concept with the template “a real, high-quality, clear and clean photo of a {$y$}” when computing the confidence scores for logit-based selectors. We found this to improve scores slightly as opposed to using the default template in the original work [39]. We attribute this to the hypothesis that CLIP should produce lower confidence scores when faced with covariate-shifted inputs, such as sketches or corrupted images, as the error rate on covariate shifts are much higher [11] than on Im-1K, which are generally clear and well-lit photographs.

### v2

@@ Appendix B Proofs
is strictly monotone increasing. The first derivative,
@@
is positive everywhere in the domain and thus, $h$ is strictly increasing on $0,\infty)$. Thus we have shown that MSP is a monotone transform of the likelihood ratio $p_{c}/p_{w}$, and Corollary 1 tells us that it is Neyman–Pearson optimal.
@@
where $L=1-d^{(1)}-d^{(2)}\geq 0$. In the binary classification case $L=0$, thus $s_{\text{RLog}}({\bm{x}})=\log\frac{p_{c}({\bm{x}})}{p_{w}({\bm{x}})}$ exactly. In the multiclass case, the $1+L/d^{(2)}$ factor varies across samples and may alter the ordering of scores. If we assume that most of the probability mass is in the top logits as stated in the theorem, which is empirically supported by high top-5 classification accuracies in prior works (Liu et al., 2021; Brock et al., 2021), then $L\approx 0$ and $s_{\text{RLog}}({\bm{x}})\approx\log\frac{p_{c}({\bm{x}})}{p_{w}({\bm{x}})}$. Since logarithm is strictly monotonically increasing, RLog is Neyman–Pearson optimal by Corollary 1. ∎
@@
The empirical likelihood of the KNN density estimator (Silverman, 2018; Zhao and Lai, 2022) is given by
@@
where $k\geq 2$, $u_{k}({\bm{z}})$ and $v_{k}({\bm{z}})$ are the Euclidean distances from $\bm{z}$} to its $k$-th nearest neighbor from $A_{c}$ and $A_{w}$ and $V_{d}$ is the unit-ball volume in $\mathbb{R}^{d}$. A classic result of non-parametric nearest neighbor density estimation (Loftsgaarden and Quesenberry, 1965) states that as $k\rightarrow\infty$ but $k/N_{c}\rightarrow 0$, $k/N_{w}\rightarrow 0$, then $\hat{p}_{c}({\bm{z}})\rightarrow p_{c}({\bm{z}})$ and $\hat{p}_{w}({\bm{z}})\rightarrow p_{w}({\bm{z}})$ for every $\bm{z}$}. In other words, the empirical KNN density estimator converges to the true density under the stated asymptotic conditions.
@@
$s_{\Delta\text{-KNN}}({\bm{z}})$ $\triangleq-\log u_{k}({\bm{z}})+\log v_{k}({\bm{z}})$
$=\frac{1}{d}\log\frac{\hat{p}_{c}({\bm{z}})}{\hat{p}_{w}({\bm{z}})}-\frac{1}{d}\log\frac{N_{w}}{N_{c}}.$ (12)
@@
Since each score is already a strictly monotone transform of $L({\bm{x}})$, we are free to re-express the scores in any other convenient monotone scale without affecting relative ordering and thus Neyman–Pearson optimality. Without loss of generality, we will let $s_{i}({\bm{x}})=\log\frac{p_{c}^{(i)}({\bm{x}})}{p_{w}^{(i)}({\bm{x}})},i=1,2$, which are identical to the original scores in terms of sample acceptance and rejection patterns. Given $\lambda\in\mathbb{R}$,
@@
Recall $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ represent the hypotheses that the classifier will make a correct and wrong prediction respectively. Assuming that the density of $\mathcal{H}_{0}$ takes the form of a tilted likelihood $p_{c}^{(1)}(p_{c}^{(2)})^{\lambda}/Z_{c}$, where $Z_{c}$ is a normalizing constant, and vice-versa for $\mathcal{H}_{1}$, then $t({\bm{x}})$ is Neyman–Pearson optimal by Corollary 1.
Appendix C Average Top $k$ $\Delta$-KNN Modification
@@
Note that $U_{i}\in[0,1]$ for all $i$. Since the joint distribution of $U_{i}$ depends only on $k$, we know from the $i$-th order statistics of $k$ i.i.d. $\mathrm{Uniform}(0,1)$ variables that each $U_{i}$ is Beta-distributed, $U_{i}\sim\mathrm{Beta}(i,k-i+1),\ \ \ 0\leq U_{1}\leq\cdots\leq U_{k}=1.$ With some algebra, $u_{i}({\bm{z}})=\frac{1}{d}\log U_{i}+\log u_{k}({\bm{z}})$. Then,
@@
as $k\to\infty$ as it is a sum of $k$ i.i.d. $\mathrm{Uniform}(0,1)$ random variables. In other words, in the asymptotic limit the average log-distance is a monotone transform of the log-distance itself. By substituting Eq. 19 back into Eq. B for distances to both correct and wrong sets, the modified $\Delta$-KNN formulation remains a monotone transform o $p_{c}/p_{w}$, thus suggesting Neyman-Pearson optimality under Corollary 1.
Appendix D Algorithm Pseudocode for Proposed Scores
Pseudocode for Algorithms 1 ($\Delta$-MDS and its linear combination) and 2 (Scoring with $\Delta$-KNN and its linear combination) are shown on the next page.
@@ Appendix D Algorithm Pseudocode for Proposed Scores
3:Initialize $\mathcal{A}_{c}^{(i)}\leftarrow\emptyset$ and $\mathcal{A}_{w}^{(i)}\leftarrow\emptyset$ for $i=1$ to $K$ $\triangleright$ Class-wise correct and incorrect feature sets
@@
3:Initialize $\mathcal{A}_{c}\leftarrow\emptyset$, $\mathcal{A}_{w}\leftarrow\emptyset$ $\triangleright$ Global sets of correct and incorrect features
@@
Appendix E Experimental Details
Image Experiments.
This section outlines the models and datasets used in the vision experiments in Section 5. For classifiers, we use the ViT-H/14 variant for DFN CLIP and the “Giant” variant of EVA, both with patch size 14. Pretrained weights are obtained from the OpenCLIP [^1] https://github.com/mlfoundations/open_clip (Cherti et al., 2023) and timm [^2] https://github.com/huggingface/pytorch-image-models libraries, respectively.
For logit-based scores with CLIP, we are required to construct logits over class concepts by taking the dot product between the text embedding $T_{\theta}({\bm{c}})$ of class concepts $\bm{c}$} and the image embedding, $\phi({\bm{x}})$, where $T_{\theta}$ is the text encoder of CLIP. Given a class label $y$, we construct the class concept with the template “a real, high-quality, clear and clean photo of a {yy}” when computing the confidence scores for logit-based selectors. We found this to improve scores slightly as opposed to using the default template in the original work (Radford et al., 2021). We attribute this to the hypothesis that CLIP should produce lower confidence scores when faced with covariate-shifted inputs, such as sketches or corrupted images, as the error rate on covariate shifts are much higher (Heng and Soh, 2025) than on Im-1K, which are generally clear and well-lit photographs.
@@ Appendix E Experimental Details
Appendix F Additional Results
The following provides additional experimental results showing risk-coverage trade-offs, hyperparameter sensitivity, and performance results on both semantic and covariance shift.
Figure 2: Hyperparameter sensitivity plots for $\lambda$ and $k$ for $\Delta$-KNN-RLog with DFN on ImageNet1K. The results for this combination are not highly sensitive to $\lambda$, while for $k$ results plateau at around $k=25$.
Figure 3: Risk-coverage curves of various selector methods for CLIP (top row) and EVA (bottom row). Our proposed methods consistently achieve the best risk-coverage tradeoff and remain stable at low coverage levels.
Table 7: Additional results under covariate shift using ResNet50 trained on ImageNet1K from the official PyTorch repository. Consistent with the findings on larger models in the main text, our proposed methods outperform all baselines. These results affirm that the gains stem from algorithmic improvements in selective classification, rather than from large-scale pretraining potentially mitigating distribution shifts.
Im-1K Im-V2 Im-S Im-C Avg (1K)
Method A N A N A N A N A N
MSP 9.67 0.248 10.32 0.257 23.4 0.196 22.6 0.203 16.5 0.226
MaxLogit 12.7 0.379 13.5 0.394 25.3 0.248 24.5 0.260 19.0 0.320
Energy 16.3 0.539 17.6 0.567 28.1 0.328 27.8 0.358 22.5 0.448
MDS 21.2 0.752 21.6 0.736 58.4 1.19 58.3 1.23 39.9 0.977
KNN 28.8 1.08 30.4 1.11 40.2 0.673 46.2 0.900 36.4 0.940
RLog 8.61 0.201 9.14 0.207 24.5 0.225 22.4 0.193 16.2 0.207
SIRC 18.9 0.648 19.6 0.652 46.2 0.844 40.5 0.729 31.3 0.718
$\Delta$-MDS 13.0 0.392 13.7 0.399 30.3 0.390 27.3 0.338 21.1 0.380
$\Delta$-KNN 15.3 0.493 16.2 0.509 36.0 0.552 33.9 0.529 25.3 0.521
$\Delta$-MDS-RLog 8.33 0.189 8.86 0.195 24.1 0.215 21.9 0.180 15.8 0.195
$\Delta$-KNN-RLog 8.48 0.195 9.02 0.202 24.3 0.221 22.2 0.189 16.0 0.202
Table 8: Additional results on semantic shift datasets, namely ImageNet-O, iNaturalist, SUN and Places. The ID distribution is ImageNet-1K.
ImageNet-O iNaturalist SUN Places Avg
A N A N A N A N A N
[…]
MSP 9.70 0.459 11.7 0.273 12.2 0.293 12.8 0.319 11.6 0.336
MDS 11.8 0.584 14.1 0.370 14.8 0.398 15.2 0.414 14.0 0.442
MaxLogit 9.75 0.462 11.9 0.282 12.6 0.311 13.4 0.343 11.9 0.350
Energy 17.7 0.930 28.1 0.935 33.2 1.14 32.9 1.13 28.0 1.03
KNN 11.3 0.552 13.2 0.334 13.9 0.360 13.9 0.362 13.1 0.402
RLog 6.21 0.253 10.0 0.205 11.2 0.254 11.5 0.267 9.73 0.245
SIRC 18.2 0.958 32.7 1.12 30.2 1.02 32.3 1.11 28.4 1.05
$\Delta$-KNN 5.93 0.237 9.42 0.181 10.6 0.228 10.4 0.220 9.09 0.217
$\Delta$-KNN-RLog 5.20 0.194 8.60 0.148 9.69 0.192 9.76 0.195 8.31 0.182
[…]
MSP 4.94 0.285 8.46 0.214 9.29 0.251 10.1 0.289 8.20 0.260
MDS 4.58 0.258 6.64 0.133 7.41 0.167 7.78 0.184 6.60 0.186
MaxLogit 6.37 0.392 10.8 0.319 11.5 0.349 12.6 0.399 10.3 0.365
Energy 8.80 0.572 14.5 0.487 14.5 0.484 15.8 0.545 13.4 0.522
KNN 4.63 0.262 6.70 0.135 7.43 0.168 7.70 0.180 6.62 0.186
RLog 3.82 0.201 6.92 0.145 8.25 0.205 8.21 0.203 6.80 0.189
SIRC 18.2 0.958 32.7 1.12 30.2 1.02 32.3 1.11 28.4 1.05
$\Delta$-MDS 3.49 0.177 5.63 0.087 6.72 0.136 6.95 0.146 5.70 0.137
$\Delta$-MDS-RLog 3.37 0.202 5.64 0.130 6.90 0.187 7.08 0.186 5.75 0.176